OpenStudy (anonymous):

How do I solve this by completing the square: x2 + 8x = -3.

OpenStudy (anonymous):

\[ (x+d)^2 = x^2+2dx+d^2 \\\ 2d = 8 \implies d = 4\implies d^2=16\\ (x+4)^2-16 = -3\\ (x+4)^2=13 \]

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

Do I win a million dollars?

OpenStudy (anonymous):

I did the same thing but it's not in the choices :/

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

A. -3 and -5 B. -4 - √13 and -4 + √13 C. -4 - √19 and -4 + √19 D. 4 - √13 and 4 + √13

OpenStudy (anonymous):

\[\begin{array}{ccr} (x+4)^2&=&13\\ |x+4| &=& \sqrt{13}\\ x+4 &=& \pm \sqrt{13} \\ x &=& -4 \pm\sqrt{13} \end{array} \]

OpenStudy (anonymous):

I salute you sir